# On an approach to canonicalizing elliptic Feynman integrals

### v2

@@ 3.1.1 Definition
The first $4$ integrals are in the top sector, and we have $7$ master integrals in total. However, if we consider the case with two unequal masses, say, $m_{3}=m_{1}\neq m_{2}$. Then, $I_{11200}=I_{21100}$ and $I_{01100}=I_{11000}$ according to symmetric relations. Thus, we only have $3$ integrals in the top sector and $5$ master integrals in total. We will study the family in $d=2-2\varepsilon$, and only focus on the top sector since the non-elliptic sectors are trivial.
@@
We first convert to the Legendre norm form and end up with the following variables by a Möbius transformation:
$\lambda=\frac{4z(x^{2}-y^{2})}{[x^{2}-(z+1)^{2}][y^{2}-(z-1)^{2}]}\,,\quad e_{5}=\frac{4x^{2}z}{(z-1)^{2}[(z+1)^{2}-x^{2}]}\,,\quad e_{6}=\frac{4z}{(z+1)^{2}-x^{2}}\,.$ (64)
For the two unequal masses case, $z$ degenerates to $1$ and the branch points become
$c_{1}=0\,,\quad c_{2}=x^{2}\,,\quad c_{3}=y^{2}\,,\quad c_{4}=4\,.$ (65)
The variables for the corresponding Legendre family are
$\lambda=\frac{4(x^{2}-y^{2})}{(x^{2}-4)y^{2}}\,,\quad e=-\frac{4}{x^{2}-4}\,.$ (66)
We consider the region $20<x<y<$ 2, $1<z<1+x$, and thus we have the order $c_{5}<c_{1}<c_{2}<c_{3}<c_{4}$. For concreteness, we consider the limit $x\to y^{-}$ where $\lambda\to 0$, and we only need the periods with subscript “00” in this limit.
3.1.2 Two unequal masses
We first consider the simpler case with two unequal masses. Following the construction in Sec. 2.2, the pre-canonical basis integrals are given by
$\phi_{1}^{\rm sr2}$ $=y\sqrt{4-x^{2}}\frac{{\rm d}z_{4}}{\sqrt{z_{4}(z_{4}-4)(z_{4}-x^{2})(z_{4}-y^{2})}}\,,$ (67)
$\phi_{2}^{\rm sr2}$ $=-z\frac{{\rm d}z_{4}}{\sqrt{z_{4}(z_{4}-4)(z_{4}-x^{2})(z_{4}-y^{2})}}\,,$ (68)
$\phi_{3}^{\rm sr2}$ $=\frac{4y(z-x^{2})}{z\sqrt{4-x^{2}}}\frac{{\rm d}z_{4}}{\sqrt{z_{4}(z_{4}-4)(z_{4}-x^{2})(z_{4}-y^{2})}}\,.$ (69)
Next, we need to transform the pre-canonical basis to the canonical basis. The transformation ${\cal T}_{\rm can}$ is given in Sec. 2.3. The explicit form of the transformation can be obtained by plugging in the variables in Eq. (64). Applying the transformation, we find the corresponding connection matrix is automatically canonical and no further transformation is needed.
3.1.3 Three unequal masses
Following our procedure, we claim the case with three unequal masses is not substantially different from the previous case. The pre-canonical basis integrals are given by
$\phi_{1}^{\rm sr3}$ $=\sqrt{[(z+1)^{2}-x^{2}][y^{2}-(z-1)^{2}]}\frac{{\rm d}z_{4}}{\sqrt{[z_{4}-(z-1)^{2}][z_{4}-(z+1)^{2}](z_{4}-x^{2})(z_{4}-y^{2})}}\,,$ (70)
$\phi_{2}^{\rm sr3}$ $=\frac{xy(z+1)[(z-1)^{2}-z_{4}]}{z_{4}(z-1)}\frac{{\rm d}z_{4}}{\sqrt{[z_{4}-(z-1)^{2}][z_{4}-(z+1)^{2}](z_{4}-x^{2})(z_{4}-y^{2})}}\,,$ (71)
$\phi_{3}^{\rm sr3}$ $=[(z-1)^{2}-z_{4}]\frac{{\rm d}z_{4}}{\sqrt{[z_{4}-(z-1)^{2}][z_{4}-(z+1)^{2}](z_{4}-x^{2})(z_{4}-y^{2})}}\,,$ (72)
$\phi_{4}^{\rm sr3}$ $=\frac{4z(x^{2}-z_{4})}{(z-1)^{2}-z_{4}}\sqrt{\frac{(y-z+1)(y+z-1)}{(z+1)^{2}-x^{2}}}\frac{{\rm d}z_{4}}{\sqrt{[z_{4}-(z-1)^{2}][z_{4}-(z+1)^{2}](z_{4}-x^{2})(z_{4}-y^{2})}}\,.$ (73)
Similarly, we only need to apply the transformation to the canonical basis, and the mixing is $\varepsilon$-factorized automatically.
3.2 Non-planar double box
3.2.1 Definition
@@
Figure 2: The non-planar double box diagram is shown in the left panel. The sub-sectors are obtained by pinching the propagators. A special case is the elliptic triangle sector, shown in the right panel. This sub-sector and the top sector are the only sectors that are elliptic, and they share the same elliptic curve. In the plot, thick lines are massive while all the others are massless, and we use the shorthand $p_{1,2,3}=p_{1}+p_{2}+p_{3}$ and $k_{1,2}=k_{1}+k_{2}$.
@@
$D_{1}$ $=k_{1}^{2}-m^{2}\,,\quad$ $D_{2}$ $=(k_{1}+p_{3})^{2}-m^{2}\,,$ (74)
$D_{3}$ $=(k_{1}+k_{2}+p_{1}+p_{2}+p_{3})^{2}-m^{2}\,,\quad$ $D_{4}$ $=(k_{1}+k_{2})^{2}-m^{2}\,,$
$D_{5}$ $=k_{2}^{2}\,,\quad$ $D_{6}$ $=(k_{2}+p_{1})^{2}\,,$
$D_{7}$ $=(k_{2}+p_{1}+p_{2})^{2}\,,\quad$ $D_{8}$ $=(k_{1}-p_{1})^{2}-m^{2}\,,$
$D_{9}$ $=(k_{2}+p_{1}+p_{3})^{2}\,.$
We have the same propagators for the electroweak double box family, except $D_{1}$ and $D_{3}$ are massless. The integral family depends on the mass $m$ and two Mandelstam variables associated with the incoming momenta $p_{1}$, $p_{2}$ and $p_{3}$:
@@
The first 4 master integrals are in the top sector. They are chosen such that their differential equations have relatively simple dependencies on the (canonical) sub-sector integrals. This will be helpful in transforming these dependencies into the canonical form. The sub-sector integrals in the elliptic triangle family themselves are well-studied in the literature Jiang:2023jmk ; Gorges:2023zgv , where canonical bases have been constructed. For remaining sub-sector integrals, we construct the canonical basis integrals with ${\rm d}\log$-forms following Chen:2020uyk ; Chen:2022lzr , and the explicit expressions for them are in App. C. Alternative canonical bases for sub-sectors are also given in the literature, such as Ahmed:2024tsg ; Becchetti:2025rrz .
Here we have two sectors sharing the same elliptic curve in the family, and we find it enlightening to study the two elliptic sectors together by applying the elliptic (triangle) cut instead of the maximal cut of the top sector, under which only the two elliptic sectors survive. The elliptic cut of master integral $I_{111110100}$ is given by [^10] We have defined $z=z_{9}-z_{6}+c_{5}$.
@@
where
@@
Branch points ${c_{1},c_{2},c_{3},c_{4}}$ can be expressed with kinematic variables $y_{s}$ and $y_{t}$ as [^11] We can also choose variables more carefully to make the branch points free of square roots with e.g. $k\equiv\sqrt{1-16y_{s}}$. However, since the family is less studied, we choose the simplest variables and make the discussions more general. A canonical basis is also reported recently in Becchetti:2025rrz .
@@
while branch points $c_{6}$ and $c_{7}$ also depend on variable $z$
$c_{6,7}=2y_{t}z-y_{t}-z\pm 2\sqrt{y_{t}(y_{t}-1)z(z-1)}\,,$ (81)
and $P_{2}(z_{6},z)$ is defined as
@@
If we further apply the cut with respect to $z_{6}$, $c_{6}$ and $c_{7}$ degenerated to a single marked point $c_{5}$
$\frac{1}{2i\pi}\oint_{z_{6}=0}\frac{1}{z_{6}\sqrt{P_{2}(z_{6},z)}}=-\frac{1}{z-c_{5}}\,,$ (83)
where
$c_{5}=y_{t}\,.$ (84)

### v3

@@ 3.1 Unequal-mass Sunrise
We use Kira Klappert:2020nbg to reduce the integrals in this family to a set of 7 master integrals, chosen as
@@
The first $4$ integrals are in the elliptic top sector, while the remaining $3$ sub-sector integrals are non-elliptic. For simplicity, we take $m_{3}=m_{1}\neq m_{2}$, and then $I_{11200}=I_{21100}$ and $I_{01100}=I_{11000}$ according to symmetric relations. We will study the family in $d=2-2\varepsilon$, where the non-elliptic master integrals are automatically canonical. We can then focus on the top sector.
@@
We consider the region $0<x<y<2$, $1<z<1+x$, where the ordering of the branch points is given by $c_{5}<c_{1}<c_{2}<c_{3}<c_{4}$. According to Eqs. (6) and (2.1), the branch points are mapped into the Legendre normal form as:
$e_{3}=\lambda=\frac{4z(x^{2}-y^{2})}{[x^{2}-(z+1)^{2}][y^{2}-(z-1)^{2}]}\,,\quad e_{5}=\frac{4x^{2}z}{(z-1)^{2}[(z+1)^{2}-x^{2}]}\,,\quad e_{6}=e_{\infty}=\frac{4z}{(z+1)^{2}-x^{2}}\,.$ (61)
Note that in the limit $x\to z-1$, we have $c_{1}\to c_{2}$ and $\lambda\to 1$.
To write down the canonical basis for the top sector, we just need to plug Eqs. (60) and Eqs. (61) into Eqs. (10), followed by converting the integrands into linear combinations of the master integrals $I_{11200}$, $I_{12100}$ and $I_{11100}$. Accidentally, the sub-sector dependence of these top-sector canonical basis turns out to be $\varepsilon$-factorized automatically. As we will see in later examples, the sub-sector dependence can be non-trivial in more complicated cases, which requires further manipulation.
3.2 An elliptic sub-sector of $2$-loop $t\bar{t}W$ production
The diagram for an elliptic sub-sector of $2$-loop $t\bar{t}W$ production is shown in Fig. 2. We follow the notations in Becchetti:2025qlu with slight modifications. The propagators are
$D_{1}$ $=k_{1}^{2}-m_{t}^{2}\,$$\quad$ $D_{2}$ $=(k_{1}-p_{2}-p_{3})^{2}\,$$\quad$ $D_{3}$ $=(k_{1}-p_{2}-p_{3}-p_{4})^{2}\,$, (62)
$D_{4}$ $=k_{2}^{2}-m_{t}^{2}\,$$\quad$ $D_{5}$ $=(k_{2}-p_{1})^{2}\,$$\quad$ $D_{6}$ $=(k_{1}+k_{2})^{2}\,$,
$D_{7}$ $=(k_{2}+p_{2}+p_{3}+p_{4})^{2}\,$$\quad$ $D_{8}$ $=(k_{1}+p_{1})^{2}\,$$\quad$ $D_{9}$ $=(k_{2}+p_{2}+p_{3})^{2}-m_{t}^{2}\,.$
The kinematic variables are defined as
$s_{45}\equiv(p_{1}+p_{2}+p_{3})^{2}\,,\quad s_{14}\equiv(p_{1}+p_{4})^{2}\,,\quad s_{23}\equiv(p_{2}+p_{3})^{2}\,.$ (63)
The scale $\mu^{2}$ is chosen as $m_{t}^{2}$, and the dimensionless variables are
$x\equiv-\frac{s_{45}}{m_{t}^{2}}\,,\quad y\equiv-\frac{s_{14}}{m_{t}^{2}}\,,\quad z\equiv-\frac{s_{23}}{m_{t}^{2}}\,,\quad k\equiv-\frac{m_{W}^{2}}{m_{t}^{2}}\,.$ (64)
Figure 2: A family contributing to 22-loop $t\bar{t}W$ production Becchetti:2025qlu is shown in the left panel. There are two four-point elliptic sectors, one of which has 4 scales and its canonical basis has been given in Becchetti:2025oyb . The remaining elliptic sector is shown in the right panel, which is more complicated with 5 scales. In the diagrams, thick orange lines represent top quarks with mass $m_{t}$, the curvy lines represent $W$ bosons with mass $m_{W}$, and the thick blue line is an off-shell particle coming from pinching while all the others are massless.
For this family, we will only consider the elliptic sector in the right panel of Fig. 2, and the starting master integrals are chosen as
$\big\{$ $I_{111111000},I_{1111110(-1)0},I_{1111110(-2)0}\big\}\,.$ (65)
The maximal cut of the master integral $I_{111111000}$ is given by
${\rm MaxCut}[I_{111111000}]=\int_{\cal C}\,u(z_{8})\frac{{\rm d}z_{8}}{\sqrt{(x+y+1)^{2}-4kz}}\,.$ (66)
The twist $u(z_{8})$ is defined by
$u(z_{8})=[P_{4}(z_{8})]^{-1/2}\prod_{i=1}^{4}(z_{8}-c_{i})^{-\beta_{i}\varepsilon}\,$, (67)
and the branch points can be expressed with the dimensionless variables as
$c_{1}=0\,,\quad c_{2}=\frac{a-b}{(x+y+1)^{2}-4kz}\,,\quad c_{3}=\frac{a+b}{(x+y+1)^{2}-4kz}\,,\quad c_{4}=4\,$, (68)
where
$a$ $=-2k^{2}z+k[x^{2}+x(y+2z+1)+2(z+1)(y-z)-x(x+y+1)(x+y-z)]\,$, (69a)
$b$ $=2\sqrt{-k[-yz(k+x)-k(x+1)y+(k+1)z(k-x)+kz^{2}+xy(x+y+1)][-z(k+z+1)+(x+y)(z+1)]}\,.$ (69b)
As before, the relevant variable changes can be obtained combining Eqs. (68), (69), (6) and (2.1). The canonical basis can then be derived by plugging the variables into Eq. (10). Due to the complicated kinematics in this family, the explicit expressions are rather lengthy and not too illuminating. We leave them in the ancillary files attached to this paper.
3.3 An elliptic branch of $2$-loop $W$-pair production via light quark-antiquark annihilation
Figure 3: The ${\cal T}_{3F}$ branch of $2$-loop $W$-pair production via light quark-antiquark annihilation He:2024iqg is shown in the left panel. The sector is reducible, and a special sub-sector is shown in the right panel. This sub-sector is the only elliptic sector. In the diagrams, thick orange lines represent top quarks with mass $m_{t}$ and the curvy lines represent $W$ bosons with mass $m_{W}$ while all the others are massless.
The diagram for the ${\cal T}_{3F}$ branch of $2$-loop $W$-pair production via light quark-antiquark annihilation is shown in Fig. 3. We follow the notations in He:2024iqg , and the propagators are
$D_{1}$ $=k_{1}^{2}\,$$\quad$ $D_{2}$ $=(k_{1}+p_{1})^{2}\,$$\quad$ $D_{3}$ $=(k_{1}-p_{2})^{2}\,$, (70)
$D_{4}$ $=k_{2}^{2}-m_{t}^{2}\,$$\quad$ $D_{5}$ $=(k_{2}+p_{1})^{2}-m_{t}^{2}\,$$\quad$ $D_{6}$ $=(k_{2}-p_{2})^{2}-m_{t}^{2}\,$,
$D_{7}$ $=(k_{1}-k_{2})^{2}-m_{t}^{2}\,$$\quad$ $D_{8}$ $=(k_{1}-p_{2}-p_{3})^{2}\,$$\quad$ $D_{9}$ $=(k_{2}-p_{2}-p_{3})^{2}\,.$
The kinematic variables are defined as
$s\equiv(p_{1}+p_{2})^{2}\,,\quad t\equiv(p_{2}+p_{3})^{2}\,.$ (71)
The scale $\mu^{2}$ is chosen as $m_{t}^{2}$, then the dimensionless variables are
$x\equiv-\frac{s}{m_{t}^{2}}\,,\quad y\equiv-\frac{t}{m_{t}^{2}}\,,\quad z\equiv-\frac{m_{W}^{2}}{m_{t}^{2}}\,.$ (72)
For this family, we are going to derive the full canonical basis without any cuts. The starting master integrals are chosen as
$\big\{$ $I_{011200110},I_{(-1)11100110},I_{(-1)11200110},I_{111100010},I_{(-1)10100110},$ (73)
$I_{010100110},I_{(-1)11100100},I_{011100100},I_{011100010},I_{100200010},$
$I_{(-1)00100110},I_{000100110},I_{010100010},I_{011100000},I_{000200200}\big\}\,.$
The first $3$ master integrals belong to the elliptic top sector. The remaining master integrals belong to non-elliptic sub-sectors, whose canonical bases can be easily constructed with $\rm d\log$}-forms.
The maximal cut of the master integral $I_{011100110}$ is given by
${\rm MaxCut}[I_{011200110}]=\int_{\cal C}\,u(z_{1})\frac{1-2\varepsilon}{\sqrt{x(x-4z)}}{\rm d}z_{1}\,$, (74)
where
$u(z_{1})=[P_{4}(z_{1})]^{-1/2}\prod_{i=1}^{4}(z_{1}-c_{i})^{-\beta_{i}\varepsilon}\,$, (75)
and the branch points can be expressed with the dimensionless variables as
$c_{1}=0\,,\quad c_{2}=\frac{2z(y-z)-xy}{x-4z}+\frac{z\sqrt{xy+(y-z)^{2}}}{x-4z}\,,\quad c_{3}=\frac{2z(y-z)-xy}{x-4z}-\frac{z\sqrt{xy+(y-z)^{2}}}{x-4z}\,,\quad c_{4}=4\,.$ (76)
After Möbius transformation, the relevant variables are
$\lambda=\frac{16z\sqrt{xy+(y-z)^{2}}}{8z\left(\sqrt{xy+(y-z)^{2}}-y+z\right)+xy(y+4)}\,,\quad e_{5}=\frac{4(x-4z)}{2z\left(-\sqrt{xy+(y-z)^{2}}-y+z-8\right)+x(y+4)}\,.$ (77)
Plugging the variables into Eq. (10), we obtained the integrands for the canonical basis under maximal cut.
At this point, it is worth pointing out that there are always ambiguities when promoting the integrands with cuts to Feynman integrals without cuts. The sub-sector dependence of the resulting integrals strongly depends on the promoting procedure. In our case, we simply convert the canonical bases in Eq. (10) into linear combinations of the top-sector master integrals in Eq. (73), without adding further components from sub-sectors. Therefore, the sub-sector dependence is determined by the choice of the top-sector master integrals. With the choice in Eq. (73), we find that the sub-sector dependence is mostly $\varepsilon$-factorized, except for $\phi_{n-1}^{\rm can}$ in Eq. (10) that involves the differential of the second kind of elliptic integral.
It has been noticed in many situations that the derivative of the top-sector integral corresponding to $\phi_{1}^{\rm can}$ leads to a simpler sub-sector dependence. In the current case, this corresponds to the derivative of
$\frac{2\varepsilon^{3}}{\varpi_{1}(\lambda)}\sqrt{-8xz\big[\sqrt{xy+(y-z)^{2}}-y+z\big]-x^{2}y(y+1)}\,I_{011200110}\,.$ (78)
Note that this is NOT the same as the derivative in (29), since the latter lives under the maximal cut with no information about sub-sectors. Using this information, we can find “good” combinations of sub-sector integrals that can be added when promoting $\phi_{n-1}^{\rm can}$ to Feynman integrals without cuts. We will employ this trick for the examples in Sec. 3.4 and Sec. 3.5 as well. When this is done (where the derivative is taken with respect to $x$), we find that there are only 3 entries in the differential equations which are not $\varepsilon$-factorized.
We now simply need to rotate away the $\varepsilon^{0}$-terms in the $3$ entries to reach a fully $\varepsilon$-factorized (canonical) differential equation. This can be done via simple one-fold integrations. The explicit expressions are rather lengthy, and we do not present them here. The final canonical basis integrals are given in ancillary files.
3.4 Non-planar double box A
3.4.1 Definition and two-variable construction under the next-to-maximal cut
@@
Figure 4: The diagram for the non-planar double box family A is shown in the left panel. The sub-sectors are obtained by pinching the propagators. A special case is the elliptic triangle sector, shown in the right panel. This sub-sector and the top sector are the only sectors that are elliptic, and they share the same elliptic curve. In the diagrams, thick blue lines are massive while all the others are massless, and we use the shorthand $p_{1,2,3}=p_{1}+p_{2}+p_{3}$ and $k_{1,2}=k_{1}+k_{2}$ hereafter.
@@
$D_{1}$ $=k_{1}^{2}-\chi m^{2}\,$$\quad$ $D_{2}$ $=(k_{1}+p_{3})^{2}-m^{2}\,$$\quad$ $D_{3}$ $=(k_{1}+k_{2}+p_{1}+p_{2}+p_{3})^{2}-\chi m^{2}\,$, (79)
$D_{4}$ $=(k_{1}+k_{2})^{2}-m^{2}\,$$\quad$ $D_{5}$ $=k_{2}^{2}\,$$\quad$ $D_{6}$ $=(k_{2}+p_{1})^{2}\,$,
$D_{7}$ $=(k_{2}+p_{1}+p_{2})^{2}\,$$\quad$ $D_{8}$ $=(k_{1}-p_{1})^{2}-m^{2}\,$$\quad$ $D_{9}$ $=(k_{2}+p_{1}+p_{3})^{2}\,$,
where we have $\chi=1$ in this example. The integral family depends on the mass $m$ and two Mandelstam variables associated with the incoming momenta $p_{1}$, $p_{2}$ and $p_{3}$:
@@
The first 4 master integrals are in the top sector. They are chosen such that their differential equations have relatively simple dependencies on the (canonical) sub-sector integrals. This will be helpful in transforming these dependencies into the canonical form. The last two master integrals in the first line of Eq. (82) belong to the elliptic triangle family, whose canonical bases are well-studied in the literature Jiang:2023jmk ; Gorges:2023zgv . For the remaining sub-sectors, we construct the canonical basis integrals with ${\rm d}\log$-forms following Chen:2020uyk ; Chen:2022lzr , and the explicit expressions for them are given in App. C. The canonical bases for these non-elliptic sub-sectors were also studied in Ahmed:2024tsg ; Becchetti:2025rrz .
In this family, we have two sectors sharing the same elliptic curve. Therefore, although a canonical basis for this family has been recently obtained in Becchetti:2025rrz , we find it enlightening to study the two elliptic sectors together by applying the next-to-maximal elliptic triangle cut instead of the maximal cut. Under the triangle cut, we have two integration variables remaining. For example, the master integral $I_{111110100}$ in the elliptic triangle sector can be written as
@@
where we have defined $z\equiv z_{9}-z_{6}+c_{5}$, and the twist is given by
@@
The branch points $c_{1},c_{2},c_{3},c_{4}$} can be expressed with kinematic variables $y_{s}$ and $y_{t}$ as
@@
while the branch points $c_{6}$ and $c_{7}$ depend also on the variable $z$:
$c_{6,7}=2y_{t}z-y_{t}-z\pm 2\sqrt{y_{t}(y_{t}-1)z(z-1)}\,.$ (86)
The polynomial $P_{2}(z_{6},z)$ is given by
@@
Since the variable $z_{6}$ appears quadratically under the square root in the twist, we can construct a $\mathrm{d}\log$ with respect to it, and arrive at
${\rm EllCut}[I_{111110100}]=\int_{\cal C}\,\bar{u}(z_{6},z)\frac{{\rm d}z_{6}}{\sqrt{P_{2}(z_{6},z)}}\wedge\frac{{\rm d}z}{\sqrt{P_{4}(z)}}=-\int_{\cal C}\,\bar{u}(z_{6},z){\rm d}\log\left(\frac{1+\sqrt{\frac{z_{6}-c_{6}}{z_{6}-c_{7}}}}{1-\sqrt{\frac{z_{6}-c_{6}}{z_{6}-c_{7}}}}\right)\wedge\frac{{\rm d}z}{\sqrt{P_{4}(z)}}\,$, ($8$)
where $\bar{u}(z_{6},z)$ is the reduced twist
Similarly, the top-sector master integral $I_{111111100}$ under the triangle cut can be written as
$=\int_{\cal C}\,\bar{u}(z_{6},z){\rm d}\log\left(\frac{1+\sqrt{\frac{c_{7}(z_{6}-c_{6})}{c_{6}(z_{6}-c_{7})}}}{1-\sqrt{\frac{c_{7}(z_{6}-c_{6})}{c_{6}(z_{6}-c_{7})}}}\right)\wedge\frac{{\rm d}z}{(z-c_{5})\sqrt{P_{4}(z)}}\,$, (91)
where the branch point $c_{5}=y_{t}$ for $z$ appears after constructing the $\mathrm{d}\log$ for $z_{6}$:
$P_{2}(0,z)=c_{6}c_{7}=(z-c_{5})^{2}\,.$ (92)
From Eqs. (88) and (90), we can see explicitly that they share the same elliptic curve defined by the degree-4 polynomial $P_{4}(z)$. Therefore, we may utilize the construction in Sec. 2.1 to write down the integrands for $z$, such that the integrands under the next-to-maximal cut have the form ${\rm d}\log\wedge\phi^{\rm can}$. Applying the Möbius transformation in Sec. 2.1, we arrive at the following variables:
Note that $\lambda$ is related to the variable $k$ in Jiang:2023jmk as
$\lambda=1-k^{2}\,,\quad k=\frac{1-\sqrt{1-16y_{s}}}{1+\sqrt{1+16y_{s}}}\,.$ (94)
We now just need to plugging the relevant variables into Eq. (10) to obtain $\phi^{\rm can}$. At this point, we would like to point out an interesting phenomenon. The integrands in Eq. (10) involve a complete elliptic integral of the third kind depending on $e_{6}$, for both the top sector and the elliptic triangle sector. However, from the results of Jiang:2023jmk , we have already known that no such integral is needed for the triangle sector. This hints at relations between complete elliptic integrals of the first kind and the third kind. [^8] One may derive similar relations from the sunrise families. In both cases, we find that such relations are associated with the fact that the integrands of the first kind and the third kind can be combined into a ${\rm d}\log$-form. In the current case, we have
$\vartheta_{0}(e_{6},\lambda)=-\frac{\varpi_{0}(\lambda)}{1+\sqrt{1-16y_{s}}}\,,\quad\vartheta_{1}(e_{6},\lambda)=-\frac{\varpi_{1}(\lambda)}{1+\sqrt{1-16y_{s}}}-1\,.$ (95)
Inserting the explicit forms of $\varpi_{i}$ and $\vartheta_{i}$, we arrive at the relations
$\Pi(1-k,1-k^{2})=\frac{(1+k)K(1-k^{2})}{2k}\,,\quad\Pi(k,k^{2})=\frac{1}{2}K(k^{2})+\frac{\pi}{4(1-k)}\,.$ (96)
The two-variable integrands encode more information than the univariate integrands for each of the two elliptic sectors. Therefore, it is interesting to investigate the differential equations of the constructed basis, especially the mixing between these two sectors. We find that the differential equations are $\varepsilon$-factorized within each of the two sectors, as expected. However, the mixing terms are not completed factorized. The problem is again associated with $\phi_{n-1}^{\rm can}$ in the top-sector. We employ the derivative trick as in Sec. 3.3 (this time with respect to $y_{s}$) to find suitable combinations of sub-sector integrals to be added. When this is done, we find that all non-$\varepsilon$-factorized terms are removed. This is of course an accident, since we have seen in Sec. 3.3 that the derivative trick is not always able to get rid of all $\varepsilon^{0}$ terms. We will see similar behaviors in the non-planar double-box family with a different mass configuration in Sec. 3.5.
3.4.2 Mixing with non-elliptic sectors
With the two elliptic sectors sorted out, we now go on to derive the full canonical basis without cuts. For that we only need to clean the mixing between the elliptic top-sector and the non-elliptic sub-sectors. We find that such mixing in this family is considerably more complicated than the previous examples, even after applying the derivative trick. Part of the reason is that elliptic integrals appear at multiple places in the canonical integrands (10). Therefore, we find it useful to study the sub-sector dependence of the pre-canonical basis (48) (promoted to Feynman integrals without cuts), before rotating it to the canonical one. In particular, if we manage to make the sub-sector dependence $\varepsilon$-factorized for the pre-canonical integrals of the first and the third kind, the same structure will be preserved when rotating to the canonical ones. Afterwards, we only need to further deal with the canonical integral of the second kind.
The coefficient matrix $\bm{A}$ in the differential equations for the pre-canonical basis $\vec{I}^{\text{pre}}$ within the top-sector has the following structure: